I need some help badley, find the vertex, the x-intercepts, and the y-intercept of the quadratic function.(remember that the x-coordinate of the vertex is -b/2a) problem y=x2-2x-3 then it says

Vertex_________
x-intercepts__________and __________
y-intercept______________
thanks for your help

Answer by ankor@dixie-net.com(15746)

I need some help badly, find the vertex, the x-intercepts, and the y-intercept of the quadratic function.(remember that the x-coordinate of the vertex is -b/(2a) problem y=x2-2x-3 then it says
:
y = x^2 - 2x -3
:
Vertex_________
Find the axis of symmetry using x = -b/(2a), in this equation: a=1, b=-2
x = -(-2)/(2*1)
x = +2/2
x = +1;
Find the vertex by finding y: substitute 1 for x in the original equation:
y = 2^1 + -2(1) - 3
y = 1 - 2 - 3
y = -4
Vertex coordinates: x=1, y=-4
:
x-intercepts__________and __________
x intercepts occur when y = 0
x^2 - 2x - 3 = 0
This will factor to:
(x-3)(x+1) = 0
x = +3
and
x = -1
:
y-intercept______________
y intercept occurs when x = 0
Substitute 0 for x in the original equation
y = 0^2 + 2(0) -3
y = -3
:
Did this help you?? Any questions?